# A finite element method to compute the damping rate and frequency of oscillating fluids inside microfluidic nozzles

### v2

@@ 1.2 Related Literature
The majority of efforts in characterizing liquid sloshing, however, were focused on linearizing the governing Navier-Stokes equations around an equilibrium meniscus and using finite elements to discretize a variational formulation of the resulting problem ([6], [2], [48], [31], [36], [35], and [15]); the latter approach results in a generalized matrix eigenvalue problem that is suitable for a modal analysis yielding the eigenmodes (velocity, pressure and meniscus surface deformation profiles) and corresponding eigenvalues, yielding the modes’ angular frequencies for inviscid liquids, and both their angular frequencies and damping rates for viscous liquids. These finite element-based efforts typically account for gravity, but vary in their inclusion of viscosity and surface tension, and in their consideration of arbitrary container geometries. Unlike some of the previously mentioned efforts not involving the use of finite elements, finite element-based approaches do not seem to simultaneously account for the effects of viscosity and surface tension. They are part of a broader research thrust to deploy model order reduction techniques for performing modal analyses of complex flows, which also include proper orthogonal decomposition (POD), balanced POD, and dynamic mode decomposition (DMD), among others [45].
@@ 1.3 Contributions
Unlike many aforementioned works on sloshing, we opted to ignore gravity and perform the linearization of the equations around a flat equilibrium meniscus. This choice simplified the analysis, but it is relatively straightforward to modify it and start from a curved meniscus instead.
@@ 2 Problem Setup and Governing Equations
Consider a nozzle full of an incompressible fluid occupying an open, bounded, regular, and connected domain $\Omega\subset\mathbb{R}^{3}$ with a piece-wise smooth $2-$manifold boundary $\partial\Omega$. We partition the boundary into three disjoint surface patches; namely, the top surface $\Gamma_{t}$, the wall surface $\Gamma_{w}$, and the equilibrium meniscus surface $\Gamma_{m}$, each of non-zero area, such that $\partial\Omega=\overline{\Gamma_{m}}\cup\overline{\Gamma_{w}}\cup\overline{\Gamma_{t}}$, where $\overline{\Gamma}$ stands for the closure of $\Gamma$. Additionally, we assume that $\overline{\Gamma_{m}}\cap\overline{\Gamma_{t}}=\emptyset$, and that $\Gamma_{m}$’s manifold boundary is $\overline{\Gamma_{m}}\cap\overline{\Gamma_{w}}$. We consider $\Gamma_{m}$ to be planar and, without loss of generality, adopt a Cartesian coordinate system $O_{xyz}$ in which $\Gamma_{m}$ lies in the $z=0$ plane with the unit outward normal to $\Omega$ on $\Gamma_{m}$ pointing towards the negative $z-$axis (see Fig. 1). The fluid inside the nozzle is assumed to be in contact with the (rigid) walls of the nozzle on $\Gamma_{w}$ and in contact with the atmosphere on $\Gamma_{m}$. We assume that $\overline{\Gamma_{m}}\cap\overline{\Gamma_{w}}$ forms a sharp edge, so that the resulting “meniscus” is pinned, i.e., the solid-liquid-gas contact curve remains fixed along this sharp edge. Finally, the fluid in the nozzle is in contact with more of the same fluid over the top surface $\Gamma_{t}$. For simplicity, we imposed a stress-free condition on $\Gamma_{t}$, consistent with neglecting the effect of gravity in the fluid.
@@
To find the oscillation modes, we will seek solutions to (1) according to the following ansatz: [^1] Or, equivalently, by applying a Laplace transform in time.
@@ 2.1 Weak Form of the Problem
This weak form can be more succinctly expressed through the following bilinear forms defined for any ${\bf u},{\bf v}\in{\cal V}$, $p,q\in{\cal P}$ and $\xi,\zeta\in\Xi$:
@@
When written in this way, it is clear that the weak form defines a GEP in which $\lambda$ is the eigenvalue and $(\hat{\bf u},{\hat{p}},\hat{\xi})$ is the eigenvector. Setting $\hat{\bf u}={\bf 0}$, $\hat{p}=0$, and $\hat{\xi}=0$ satisfies these equations for any $\lambda\in\mathbb{C}$. However, we are looking for nontrivial solutions, and those are obtained for selected values of $\lambda$.
Notice that both bilinear forms $g$ and $h$ are symmetric, so (10d) defines a symmetric GEP.
It is shown in 6.3 that all eigenvalues need to have a negative real part, i.e., $\mathfrak{Re}(\lambda)<0$, so the amplitude of each mode decays in time. Additionally, it is trivial to see that if $\lambda$ is an eigenvalue and $(\hat{\bf u},\hat{p},\hat{\xi})$ is its eigenvector, then its complex conjugate $\overline{\lambda}$ is also an eigenvalue with $(\overline{\hat{\bf u}},\overline{\hat{p}},\overline{\hat{\xi}})$ as its eigenvector. Therefore, oscillatory modes of the fluid come in pairs with frequencies of opposite signs (imaginary parts of $\lambda$ and $\overline{\lambda}$) and the same relaxation time (real parts of $\lambda$ and $\overline{\lambda}$). [^2] Here $\overline{\zeta}$ indicates the complex conjugate of a complex number $\zeta\in\mathbb{C}$.
In general, we will be interested in finding all modes whose eigenvalue $\lambda$ satisfy $\eta\tau=-\mathfrak{Re}(\lambda)\tau<1$, for some characteristic time $\tau>0$ we choose. This corresponds to seeking all modes for which $|\mathfrak{Re}(\lambda)|$ is smaller than $1/\tau$, i.e., which will take longer to decay than the chosen relaxation time $\tau$.
@@ 2.2 Restriction to 2D
To simplify the notations hereafter, we do not make a distinction between the 2D set $\breve{\Omega}\subset\mathbb{R}^{2}$, used in (12) or (13), and its embeddings in 3D as $\Omega\cap\{(x,0,z)\in\mathbb{R}^{3}\}$ or $\Omega\cap\{(r,0,z)\in\mathbb{R}^{3}\}$, respectively. Additionally, because $\hat{\bf u}$, $\hat{p}$, and $\hat{\xi}$ change only over $\breve{\Omega}$, we will indistinctly treat them as functions over $\Omega$ or $\breve{\Omega}$.
@@ 3 Finite Element Discretization
We discretize the 2D restriction of the problem in §2.2. We mesh $\breve{\Omega}$ with triangular elements such that $\gamma_{m}$, $\gamma_{t}$, and $\gamma_{w}$ are each meshed by a distinct collection of edges in the mesh, and such that every triangle has at least one node in the interior of $\breve{\Omega}$. [^3] This is a sufficient requirement to satisfy the inf-sup condition (see 6.4). On curved boundaries, a collection of edges in the mesh interpolates the boundary at the vertices, as customary. We did not implement an isoparametric construction to more accurately approximate the domain.
@@ 3.2 Implementation and Solution of the Matrix GEP
The matrix $G$ can display poor conditioning which leads to problems with the convergence of some components of the eigenvectors. To address this problem, we left-multiply both sides of (17b) by a diagonal preconditioner $\Lambda$ with diagonal components given by $\Lambda_{ii}=1/\max_{j=1,\ldots,n}|G_{ij}|$ for $i=1,\ldots,n$, where $n$ is the dimension of the GEP (17b). This operation does not change the eigenvectors or eigenvalues of (17b), but it does transform the matrices of the system into non-symmetric ones. It is possible to keep the symmetry by preconditioning from the right as well, but we did not do it here.
@@ 4.1.1 Inviscid Case
From (21), it follows that the expected dimensionless angular frequencies are $\omega_{1}=5.5683$, $\omega_{2}=15.7496$ and $\omega_{3}=28.9339$. We perform FEA simulations for a series of Reynolds numbers ranging from 1004 to 8034. Figure 4 demonstrates convergence of the angular frequencies computed with FEA and varying viscosity values to the analytical values for the inviscid case, (21), when $n=1,2,3$.
(a) Mode $n=1$
(b) Mode $n=2$
(c) Mode $n=3$
Figure 4: Comparison of the angular frequencies obtained from the FEA solutions for the three lowest-angular-frequency modes of the viscous capillary wave with their counterparts computed via (24) for (a) mode $n=1$, (b) mode $n=2$, and (c) mode $n=3$. For $\textsf{Re}^{-1}$ approaching zero, the angular frequencies from the FEA solutions converge to the analytical expressions for the inviscid capillary wave given by (21). The values computed from the FEA solutions have largely converged, as the comparison between the two levels of refinements suggests.
Finally, Fig. 5 compares the profiles of the velocity components from the FEA solutions with their analytical counterparts given by $\hat{\bf u}_{n}=\nabla\hat{\phi}_{n}$, with $\hat{\phi}_{n}$ given by (19), for $n=1$. Note that for this example $\hat{\bf u}_{1}$ can be chosen to have all real components. In contrast, the FEA solutions are computed with a small value of $\textsf{Re}^{-1}$, so it is not necessarily possible to make all components of the computed velocity field real. However, the FEA solution can be multiplied by a complex scalar to make the real part of the components much larger than the imaginary part (in this case two orders of magnitude larger), and normalized so that the analytic and numerical solution have the same maximum value for some field, in this case, the absolute value of the real part of the velocity component in the $x-$direction. We show the real part of the resulting velocities in the figure.
@@
Figure 5: Comparison of the analytical and computed values of $\hat{u}_{x}$ and $\hat{u}_{z}$ of a planar inviscid capillary wave for $n=1$. The analytical values followed from $\nabla\hat{\phi}_{n}$ for $\hat{\phi}_{n}$ in (19).
@@ 4.1.2 Viscous Case
Figure 4 compares the angular frequencies from (24) with the estimates from the FEA solutions for $n=1,2,3$. Figure 6 shows a similar comparison for the damping rates. For both quantities, we find a good agreement between the FEA estimates and the values obtained from (24), with deviations becoming larger as $\textsf{Re}^{-1}$ increases and the use of inviscid flow modes is no longer justified in deriving (24).
@@
Figure 6: Comparison of the damping rates obtained from the FEA solutions for the three lowest-angular-frequency modes of a planar viscous capillary wave with their counterparts computed via (24) for (a) mode 1, (b) mode 2, and (c) mode 3. As in Fig. 4, the values computed from the FEA solutions have largely converged, as the comparison between the two levels of refinements suggests.
@@ 6.3 Sign of the Real Part of the Eigenvalues of the Continuous Problem
We formally show next that for any non-trivial solution $\hat{\mathbf{u}},\hat{p},\hat{\xi}$ of the weak form (8), [^5] For simplicity of notation, we drop the $\hat{\cdot}$ notation in the rest of this section. the eigenvalue $\lambda$ needs to have a negative real part, i.e., $\mathfrak{Re}(\lambda)<0$. To see that $\mathfrak{Re}(\lambda)<0$, notice first that for a solution $({\bf u},p,\xi)$with ${\bf u}\not={\bf 0}$, $a({\bf u},\overline{{\bf u}})>0$, $c({\bf u},\overline{\bf u})>0$ (from (26c)), and $s(\overline{\xi},\xi)\geq 0$, while $0=b({\bf u},\overline{p})=\overline{b(\overline{\bf u},p)}$ from (10b). Therefore

### v3

@@ 1.2 Related Literature
The majority of efforts in characterizing liquid sloshing, however, were focused on linearizing the governing Navier-Stokes equations around an equilibrium meniscus and using finite elements to discretize a variational formulation of the resulting problem ([6], [2], [48], [31], [36], [35], and [15]); the latter approach results in a generalized matrix eigenvalue problem that is suitable for a modal analysis yielding the eigenmodes (velocity, pressure and meniscus surface deformation profiles) and corresponding eigenvalues, yielding the modes’ angular frequencies for inviscid liquids, and both their angular frequencies and damping rates for viscous liquids. These finite element-based efforts typically account for gravity, but do not simultaneously account for viscosity, surface tension, and arbitrary container geometries. They are part of a broader research thrust to deploy model order reduction techniques for performing modal analyses of complex flows, which also include proper orthogonal decomposition (POD), balanced POD, and dynamic mode decomposition (DMD), among others [45].
@@ 2 Problem Setup and Governing Equations
Consider a nozzle full of an incompressible fluid occupying an open, bounded, regular, and connected domain $\Omega\subset\mathbb{R}^{3}$ with a piece-wise smooth $2-$manifold boundary $\partial\Omega$. We partition the boundary into three disjoint surface patches; namely, the top surface $\Gamma_{t}$, the wall surface $\Gamma_{w}$, and the equilibrium meniscus surface $\Gamma_{m}$, each of non-zero area, such that $\partial\Omega=\overline{\Gamma_{m}}\cup\overline{\Gamma_{w}}\cup\overline{\Gamma_{t}}$, where $\overline{\Gamma}$ stands for the closure of $\Gamma$. Additionally, we assume that $\overline{\Gamma_{m}}\cap\overline{\Gamma_{t}}=\emptyset$, and that $\Gamma_{m}$’s manifold boundary is $\overline{\Gamma_{m}}\cap\overline{\Gamma_{w}}$. We consider $\Gamma_{m}$ to be planar and, without loss of generality, adopt a Cartesian coordinate system $O_{xyz}$ in which $\Gamma_{m}$ lies in the $z=0$ plane with the unit outward normal to $\Omega$ on $\Gamma_{m}$ pointing towards the negative $z-$axis (see Fig. 1). The fluid inside the nozzle is assumed to be in contact with the (rigid) walls of the nozzle on $\Gamma_{w}$ and in contact with the atmosphere on $\Gamma_{m}$. We assume that $\overline{\Gamma_{m}}\cap\overline{\Gamma_{w}}$ forms a sharp edge, so that the resulting “meniscus” is pinned, i.e., the solid-liquid-gas contact curve remains fixed along this sharp edge. Finally, the fluid in the nozzle is in contact with more of the same fluid over the top surface $\Gamma_{t}$. For simplicity, we impose a stress-free condition on $\Gamma_{t}$, consistent with neglecting the effect of gravity in the fluid.
@@
To find the oscillation modes, we seek solutions to (1) according to the following ansatz: [^1] Or, equivalently, by applying a Laplace transform in time.
@@ 2.1 Weak Form of the Problem
The weak form can be more succinctly expressed through the following bilinear forms defined for any ${\bf u},{\bf v}\in{\cal V}$, $p,q\in{\cal P}$ and $\xi,\zeta\in\Xi$:
@@
When written in this way, it is clear that the weak form defines a GEP in which $\lambda$ is the eigenvalue and $(\hat{\bf u},{\hat{p}},\hat{\xi})$ is the eigenvector. Setting $\hat{\bf u}={\bf 0}$, $\hat{p}=0$, and $\hat{\xi}=0$ satisfies these equations for any $\lambda\in\mathbb{C}$. However, we are looking for nontrivial solutions, and those are obtained for selected values of $\lambda$. Notice that both bilinear forms $g$ and $h$ are symmetric, so (10d) defines a symmetric GEP.
We show in 6.3 that all eigenvalues need to have a negative real part, i.e., $\mathfrak{Re}(\lambda)<0$, so the amplitude of each mode decays in time. Additionally, it is trivial to see that if $\lambda$ is an eigenvalue and $(\hat{\bf u},\hat{p},\hat{\xi})$ is its eigenvector, then its complex conjugate $\overline{\lambda}$ is also an eigenvalue with $(\overline{\hat{\bf u}},\overline{\hat{p}},\overline{\hat{\xi}})$ as its eigenvector. Therefore, oscillatory modes of the fluid come in pairs with frequencies of opposite signs (imaginary parts of $\lambda$ and $\overline{\lambda}$) and the same relaxation time (real parts of $\lambda$ and $\overline{\lambda}$). [^2] Here $\overline{\zeta}$ indicates the complex conjugate of a complex number $\zeta\in\mathbb{C}$.
In general, we are interested in finding all modes whose eigenvalue $\lambda$ satisfy $\eta\tau=-\mathfrak{Re}(\lambda)\tau<1$, for some characteristic time $\tau>0$ we choose. This corresponds to seeking all modes for which $|\mathfrak{Re}(\lambda)|$ is smaller than $1/\tau$, i.e., which will take longer to decay than the chosen relaxation time $\tau$.
@@ 2.2 Restriction to 2D
To simplify the notations hereafter, we do not make a distinction between the 2D set $\breve{\Omega}\subset\mathbb{R}^{2}$, used in (12) or (13), and its embeddings in 3D as $\Omega\cap\{(x,0,z)\in\mathbb{R}^{3}\}$ or $\Omega\cap\{(r,0,z)\in\mathbb{R}^{3}\}$, respectively. Additionally, because $\hat{\bf u}$, $\hat{p}$, and $\hat{\xi}$ change only over $\breve{\Omega}$, we indistinctly treat them as functions over $\Omega$ or $\breve{\Omega}$.
@@ 3 Finite Element Discretization
We discretize the 2D restriction of the problem in §2.2. We mesh $\breve{\Omega}$ with triangular elements such that $\gamma_{m}$, $\gamma_{t}$, and $\gamma_{w}$ are each meshed by a distinct collection of edges in the mesh, and such that every triangle has at least one node in the interior of $\breve{\Omega}$. [^3] This is a sufficient requirement to satisfy the inf-sup condition (see 6.4). On curved boundaries, a collection of edges in the mesh interpolates the boundary at the vertices, as customary. We do not implement an isoparametric construction to more accurately approximate the domain.
@@ 3.2 Implementation and Solution of the Matrix GEP
The matrix $G$ can display poor conditioning which leads to problems with the convergence of some components of the eigenvectors. To address this problem, we left-multiplied both sides of (17b) by a diagonal preconditioner $\Lambda$ with diagonal components given by $\Lambda_{ii}=1/\max_{j=1,\ldots,n}|G_{ij}|$ for $i=1,\ldots,n$, where $n$ is the dimension of the GEP (17b). This operation does not change the eigenvectors or eigenvalues of (17b), but it does transform the matrices of the system into non-symmetric ones. It is possible to keep the symmetry by preconditioning from the right as well, but we did not do it here.
@@ 4.1.1 Inviscid Case
Figure 4 demonstrates that the angular frequencies for modes $n=1,2,3$ obtained from FEA solutions for Reynolds numbers ranging from 251 to 8034 converge to their counterparts computed with (21) in the limit of $\textsf{Re}^{-1}$ approaching zero.
(a) Mode $n=1$ (scale)
(b) Mode $n=2$ (scale)
(c) Mode $n=3$ (scale)
(d) Mode $n=1$ (detail)
(e) Mode $n=2$ (detail)
(f) Mode $n=3$ (detail)
Figure 4: Comparison of the angular frequencies of a viscous capillary wave solution obtained with both FEA and analytically, i.e., (24). Also included are the values for an inviscid capillary wave, given by (21). The comparison is shown for the three lowest-angular-frequency modes: (a) mode $n=1$, (b) mode $n=2$, and (c) mode $n=3$. The top row provides a sense of scale, highlighting the fact that the discrepancy between the FEA computed and analytical viscous values remains small (below 0.1%) even at low Reynolds numbers. The bottom row provides a more detailed view, highlighting both the convergence of the FEA computed values to the analytical values as the Reynolds number grows, and the fact that the FEA computed values are largely mesh-converged, as a comparison between the two levels of refinement (i.e., number of finite elements) suggests. The inviscid values in red are of course constant with Reynolds number, and confirm that both viscous solutions approach them as the Reynolds number grows.
Finally, Fig. 5 compares the profiles of the velocity components from the FEA solution for a case  [^5] This case was inspired by a domain filled with liquid aluminum in contact with argon at atmospheric conditions, which is discussed in more detail in Section 4.2. with $\textsf{Re}=710$ with its analytical inviscid counterpart $\hat{\bf u}_{1}=\nabla\hat{\phi}_{1}$, where $\hat{\phi}_{n}$ is given by (19) for $n=1$. Note that for this example $\hat{\bf u}_{1}$ can be chosen to have all real components. In contrast, the FEA solution is computed with a small value of $\textsf{Re}^{-1}$, so it is not necessarily possible to make all components of the computed velocity field real. However, the FEA solution can be multiplied by a complex scalar to make the real part of the components much larger than the imaginary part (in this case two orders of magnitude larger), and normalized so that the analytical and numerical solution have the same maximum value for some field, in this case, the absolute value of the real part of the velocity component in the $x-$direction. We show the real part of the resulting velocity components in the figure.
@@
Figure 5: Comparison of the analytically and FEA computed functions $\hat{u}_{x}$ and $\hat{u}_{z}$ corresponding to the velocity field of a planar inviscid capillary wave with $n=1$. The FEA values were computed for $\textsf{Re}=710$, and the analytical values follow as $\hat{u}=\nabla\hat{\phi}_{n}$ for $\hat{\phi}_{n}$ in (19).
@@ 4.1.2 Viscous Case
Figure 4 compares the angular frequencies from (24) with the estimates from the FEA solutions with Reynolds numbers ranging from 251 to 8034 for $n=1,2,3$. Figure 6 shows a similar comparison for the damping rates. For both quantities, we find a good agreement between the FEA estimates and the values obtained from (24), with deviations becoming larger as $\textsf{Re}^{-1}$ increases and the use of inviscid flow modes is no longer justified in deriving (24).
@@
Figure 6: Comparison of the damping rates obtained from the FEA solutions for the three lowest-angular-frequency modes of a planar viscous capillary wave with their counterparts computed via (24) for (a) mode 1, (b) mode 2, and (c) mode 3. As in Fig. 4, the values computed from the FEA solutions are largely mesh-converged, as a comparison between the two levels of refinement suggests.
@@ 4.2 Mesh Convergence on a Cylindrical Domain
(a) $e_{u}^{h}$ for mode 1
(b) $e_{p}^{h}$ for mode 1
(c) $e_{\xi}^{h}$ for mode 1
(d) $e_{u}^{h}$ for mode 2
(e) $e_{p}^{h}$ for mode 2
(f) $e_{\xi}^{h}$ for mode 2
(g) $e_{u}^{h}$ for mode 3
(h) $e_{p}^{h}$ for mode 3
(i) $e_{\xi}^{h}$ for mode 3
@@ 4.3 Mesh Convergence on an Arbitrary Axisymmetric Domain
(a) $e_{u}^{h}$ for mode 1
(b) $e_{p}^{h}$ for mode 1
(c) $e_{\xi}^{h}$ for mode 1
(d) $e_{u}^{h}$ for mode 2
(e) $e_{p}^{h}$ for mode 2
(f) $e_{\xi}^{h}$ for mode 2
(g) $e_{u}^{h}$ for mode 3
(h) $e_{p}^{h}$ for mode 3
(i) $e_{\xi}^{h}$ for mode 3
@@ 6.3 Sign of the Real Part of the Eigenvalues of the Continuous Problem
We formally show next that for any non-trivial solution $\hat{\mathbf{u}},\hat{p},\hat{\xi}$ of the weak form (8), [^7] For simplicity of notation, we drop the $\hat{\cdot}$ notation in the rest of this section. the eigenvalue $\lambda$ needs to have a negative real part, i.e., $\mathfrak{Re}(\lambda)<0$.
To see that $\mathfrak{Re}(\lambda)<0$, notice first that for a solution $({\bf u},p,\xi)$with ${\bf u}\not={\bf 0}$, $a({\bf u},\overline{{\bf u}})>0$, $c({\bf u},\overline{\bf u})>0$ (from (26c)), and $s(\overline{\xi},\xi)\geq 0$, while $0=b({\bf u},\overline{p})=\overline{b(\overline{\bf u},p)}$ from (10b). Therefore